Two batteries of emf 3V and 1V and of internal resistance 1 ohm each are connected in parallel. Find the effective emf of the combination.

Grade:12th pass

2 Answers

Abhishek Singh
7 years ago
This depends on how do you connect them.
If you connect both of them with positive polarity together then e = ( (3/1)+(1/1) ) /(  (1/1) + (1/1)     )
= 2 volts
if you connect with the opposite polarity together  e = ( (3/1)  –  (1/1) ) /(  (1/1) + (1/1))
= 1 volts
this is accoring to the rule E =\frac{\sum (e/r)}{\sum (1/r)}
 
Saurabh Kumar
7 years ago
it would be, E1/r1 + E2/r2/(1/r1 + 1/r2)